# Split Covariance Intersection Filter Based Visual Localization With Accurate AprilTag Map For Warehouse Robot Navigation

### v2

@@ I INTRODUCTION
Robot localization is fundamental for mobile robotics and is involved in a large variety of practical applications [1, 2, 3, 4, 5, 6, 7]. Visual localization plays an important role in mobile robotics, thanks to its commercial competitiveness. It is true that a visual localization system may be susceptible to light conditions and has comparatively short perception range in contrast with LiDAR-based localization which also plays an important role in mobile robotics. However, these limitations of visual localization are naturally overcome in the context of indoor mobile robotics such as warehouse applications, because in indoor environments artificial light conditions are usually stable and the need for long-range perception (such as that in outdoor applications) is rare.
Visual simultaneous localization and mapping (SLAM) [8, 9, 10, 11, 12] is a special form of visual localization, usually without a priori map. Despite its popularity and its merits for flexible exploration in an unknown environment, visual SLAM without a priori map is unlikely to be a proper solution for a known environment of which an accurate map can be established. After all, an accurate map can largely facilitate visual localization.
@@
AprilTag measurements may be nonlinear partial measurements. For example, the performance of the AprilTag detection module deteriorates as the view distance and the view angle increase [21, 22]. Sometimes the AprilTag detection module may output distance measurements only instead of complete pose measurements [17]. This is likely to happen when its sub-step of homography computation encounters the view singularity problem due to visual data errors and noises. Nonlinear partial measurements can be handled by the well-known extended Kalman filter (EKF) in the spirit of local linearization [18] [23].
AprilTag measurements tend to have (unknown) temporal correlation; however, this cannot be reasonably handled by the EKF. The split covariance intersection filter (Split CIF) [24] [25] is adopted to handle temporal correlation among AprilTag measurements — The Split CIF, which may be regarded as a generalization of both the (extended) Kalman filter and the covariance intersection filter, can reasonably handle both known independent information and unknown correlated information in source data. It has been applied in a number of intelligent vehicle applications [26, 27, 28].
@@ III-A Split Covariance Intersection Filter
Where ${\omega_{opt}}\in[0,1]$, and ${\omega_{opt}}$ is determined by solving a convex optimization problem (see [23] for details). In addition, about the Split CIF, it can be regarded as a generalization of the Kalman filter, as shown in above equation. Specially, let ${{\bf{P}}_{1d}}$ and ${{\bf{P}}_{2d}}$ be zero, and above equation will become similar to the Kalman filter.
Concerning the instability of the AprilTag detection module, the quality of measurements will be different and there exists many low accurate observations that can be regarded as outliers. We incorporate a measurement adaptive mechanism to handle outliers in AprilTag measurements. For these different quality observations, we process them with different way, which also can be called soft abandon mechanism. It is assumed that the predicted value is ${{\widetilde{\bf{X}}}_{k+1/k}}$, and then the current measurements is ${\bf{Z}}_{{k+1}}$. When $||{{\bf{Z}}_{k+1}}-{{\widetilde{\bf{X}}}_{k+1/k}}||$ is greater than a screening threshold, ${\bf{Z}}_{{k+1}}$ is discarded. When observations are not discarded but are of lower quality, we drive an adaptive and accurate observation noise model which fits the relationship of observation error and view distance (denote as $L$) and view angle (denote as $\alpha$) according to experimental error statistics to dynamically evaluate the observation uncertainty by its noise covariance as follows:
${\bf{R}}_{k+1}=0.25(L/(\alpha^{2}))||{{\widetilde{\bf{X}}}_{k+1/k}}-{{\bf{Z}}_{k+1}}||$ (2)
@@ III-B1 Robot State Dynamic Initialization
Before the recursive robot state estimation, the robot can obtain its initial global pose ${{\bf{X}}_{{}_{0}}}$ by the dynamic initialization, which can be calculated according to the first frame of AprilTag detection data matching with the map. And the state covariance ${{\bf{P}}_{{}_{0}}}$, process noise covariance ${{\bf{Q}}_{0}}$ can be set according to motion and visual sensors or real applications. Specifically, the ${{\bf{Q}}}$ can be set according to product parameters of IMU and odometer that input the control variables. In addition, the initial independent covariance ${{\bf{P}}_{{}_{i,0}}}$ can be the same as ${{\bf{P}}_{{}_{0}}}$, and dependent covariance ${{\bf{P}}_{{}_{d,0}}}$ can be set to 00. And the independent part of observation noise covariance ${\bf{R}}_{{i,k+1}}$ can be replaced by the adaptive observation noise model in equation (2), and the initial dependent part can be regarded zero in the complete update observation process. Meanwhile, in partial update process, the initial independent part of split observation noise covariance can be set according to the statistic of the error of visual sensor detecting the distance in real applications, and the initial dependent part can be regarded zero.
@@ III-B4 Robot State Update with AprilTag Nonlinear Partial Measurements
In the spirit of local linearization on which the extended Kalman filter (EKF) relies [23], the Split CIF also enables update with nonlinear partial measurements. More specifically, suppose the warehouse robot has only a measurement of its distance to an AprilTag that has the globally-registered position {$x_{g},y_{g}$} in the accurate AprilTag map established a priori. The nonlinear partial measurement model is as follows, and we can locally linearize the measurement model about the prediction state ${\tilde{x}}_{k+1/k},{\tilde{y}}_{k+1/k}$} for the current period as follows:
@@ III-B5 Back-projection
In addition, our solution also consider the situation that the recursive estimation process received measurements are delayed since the image transmission and processing of the AprilTag is delayed for few seconds, which is common in practical applications and need to be resolved. The general solution that regard the error caused by delay as the random measurement error is not suitable for such large delay. Therefore, we apply the back-projection (BP) using motion data technology [38, 26] to compensate the delayed state estimate. Specifically, suppose the robot pose estimation (either via prediction or via update) at current time period and previous time periods are ${\bf X}_{k}$ and ${\bf X}_{k-1},{\bf X}_{k-2},...,{\bf X}_{k-m},...$. For example, we can treat the motion data period as the estimate time period. Besides, store previous motion data ${\bf u}_{t},{\bf u}_{k-1},...$ as well. We can store previous pose estimates and motion data in dynamically-adjusted queue (e.g. arrays) structures. Herein, the system prediction model is denoted as $g(\cdot)$, and it has {${{\bf X}_{k},{\bf P}_{k}}$} = $g({{\bf X}_{k},{\bf P}_{k}},{\bf u}_{k})$. Then, imagine at current time $k$, we receive a AprilTag measurement result whose timestamp is at (or near) time period $k-m$. In other words, we have an AprilTag measurement ${\bf Z}_{k-m}$. What the back-projection using motion data does is as follows:
a) Fuse (i.e. update) ${\bf X}_{k-m}$ with ${\bf Z}_{k-m}$ via the Split CIF as if ${\bf Z}_{k-m}$ was available at time $k-m$. Suppose the fused new result is {${{\bf X}_{k-m,new},{\bf P}_{k-m,new}}$}.
b) Back-project to update all pose estimates (together with their covariance) from time $k-m$ to $k$ in an iterative way as follows:

### v3

@@ I INTRODUCTION
Robot localization is fundamental for mobile robotics and is involved in a large variety of practical applications [1, 2, 3, 4, 5, 6, 7, 8]. Visual localization plays an important role in mobile robotics, thanks to its commercial competitiveness. It is true that a visual localization system may be susceptible to light conditions and has comparatively short perception range in contrast with LiDAR-based localization which also plays an important role in mobile robotics. However, these limitations of visual localization are naturally overcome in the context of indoor mobile robotics such as warehouse applications, because in indoor environments artificial light conditions are usually stable and the need for long-range perception (such as that in outdoor applications) is rare.
Visual simultaneous localization and mapping (SLAM) [9, 10, 11, 12, 13, 14] is a special form of visual localization, usually without a priori map. Despite its popularity and its merits for flexible exploration in an unknown environment, visual SLAM without a priori map is unlikely to be a proper solution for a known environment of which an accurate map can be established. After all, an accurate map can largely facilitate visual localization.
@@
AprilTag measurements may be nonlinear partial measurements. For example, the performance of the AprilTag detection module deteriorates as the view distance and the view angle increase [23, 24]. Sometimes the AprilTag detection module may output distance measurements only instead of complete pose measurements [19]. This is likely to happen when its sub-step of homography computation encounters the view singularity problem due to visual data errors and noises. Nonlinear partial measurements can be handled by the well-known extended Kalman filter (EKF) in the spirit of local linearization [20] [25] [26].
AprilTag measurements tend to have (unknown) temporal correlation; however, this cannot be reasonably handled by the EKF. The split covariance intersection filter (Split CIF) [27] [28] [29] is adopted to handle temporal correlation among AprilTag measurements — The Split CIF, which may be regarded as a generalization of both the (extended) Kalman filter and the covariance intersection filter, can reasonably handle both known independent information and unknown correlated information in source data. It has been applied in a number of intelligent vehicle applications [30, 31, 32, 33, 34, 35, 36, 37].
@@ III-A Split Covariance Intersection Filter
Where ${\omega_{opt}}\in[0,1]$, and $\omega_{opt}$} is determined by solving a convex optimization problem (see [25] [26] for details). In addition, about the Split CIF, it can be regarded as a generalization of the Kalman filter, as shown in above equation. Specially, let ${\bf{P}}_{1d}$} and ${\bf{P}}_{2d}$} be zero, and above equation will become similar to the Kalman filter.
Concerning the instability of the AprilTag detection module, the quality of measurements will be different and there exists many low accurate observations that can be regarded as outliers. We incorporate a measurement adaptive mechanism to handle outliers in AprilTag measurements. For these different quality observations, we process them with different way, which also can be called soft abandon mechanism. It is assumed that the predicted value is ${\widetilde{\bf{X}}}_{k+1/k}$}, and then the current measurements is ${\bf{Z}}_{{k+1}}$. When ${{\bf{Z}}_{k+1}}-{{\widetilde{\bf{X}}}_{k+1/k}}||$ is greater than a screening threshold, ${\bf{Z}}_{{k+1}}$ is discarded. When observations are not discarded but are of lower quality, we drive an adaptive and accurate observation noise model which fits the relationship of observation error and view distance (denote as $L$) and view angle (denote as $\alpha$) according to experimental error statistics to dynamically evaluate the observation uncertainty by its noise covariance as follows:
[…]
@@ III-B1 Robot State Dynamic Initialization
Before the recursive robot state estimation, the robot can obtain its initial global pose ${\bf{X}}_{{}_{0}}$} by the dynamic initialization, which can be calculated according to the first frame of AprilTag detection data matching with the map. And the state covariance ${\bf{P}}_{{}_{0}}$}, process noise covariance ${\bf{Q}}_{0}$} can be set according to motion and visual sensors or real applications. Specifically, the $\bf{Q}$}} can be set according to product parameters of IMU and odometer that input the control variables. In addition, the initial independent covariance ${\bf{P}}_{{}_{i,0}}$} can be the same as ${\bf{P}}_{{}_{0}}$}, and dependent covariance ${\bf{P}}_{{}_{d,0}}$} can be set to 0. And the independent part of observation noise covariance ${\bf{R}}_{{i,k+1}}$ can be replaced by the adaptive observation noise model in equation (2), and the initial dependent part can be regarded zero in the complete update observation process. Meanwhile, in partial update process, the initial independent part of split observation noise covariance can be set according to the statistic of the error of visual sensor detecting the distance in real applications, and the initial dependent part can be regarded zero.
@@ III-B4 Robot State Update with AprilTag Nonlinear Partial Measurements
In the spirit of local linearization on which the extended Kalman filter (EKF) relies [25] [26], the Split CIF also enables update with nonlinear partial measurements. More specifically, suppose the warehouse robot has only a measurement of its distance to an AprilTag that has the globally-registered position $x_{g},y_{g}$} in the accurate AprilTag map established a priori. The nonlinear partial measurement model is as follows, and we can locally linearize the measurement model about the prediction state ${\tilde{x}}_{k+1/k},{\tilde{y}}_{k+1/k}$} for the current period as follows:
@@ III-B5 Back-projection
In addition, our solution also consider the situation that the recursive estimation process received measurements are delayed since the image transmission and processing of the AprilTag is delayed for few seconds, which is common in practical applications and need to be resolved. The general solution that regard the error caused by delay as the random measurement error is not suitable for such large delay. Therefore, we apply the back-projection (BP) using motion data technology [47, 30] to compensate the delayed state estimate. Specifically, suppose the robot pose estimation (either via prediction or via update) at current time period and previous time periods are ${\bf X}_{k}$ and ${\bf X}_{k-1},{\bf X}_{k-2},...,{\bf X}_{k-m},$.... For example, we can treat the motion data period as the estimate time period. Besides, store previous motion data ${\bf u}_{t},{\bf u}_{k-1},$... as well. We can store previous pose estimates and motion data in dynamically-adjusted queue (e.g. arrays) structures. Herein, the system prediction model is denoted as $g(\cdot)$, and it has ${\bf X}_{k},{\bf P}_{k}$}} = $g({{\bf X}_{k},{\bf P}_{k}},{\bf u}_{k})$. Then, imagine at current time $k$, we receive a AprilTag measurement result whose timestamp is at (or near) time period kmk-m. In other words, we have an AprilTag measurement ${\bf Z}_{k-m}$. What the back-projection using motion data does is as follows:
a) Fuse (i.e. update) ${\bf X}_{k-m}$ with ${\bf Z}_{k-m}$ via the Split CIF as if ${\bf Z}_{k-m}$ was available at time kmk-m. Suppose the fused new result is ${\bf X}_{k-m,new},{\bf P}_{k-m,new}$}}.
b) Back-project to update all pose estimates (together with their covariance) from time kmk-m to $k$ in an iterative way as follows: